OpenStudy (anonymous):

Tommy is making a birdhouse at a craft store. He can choose 1 of 4 house styles, 1 of 3 types of wood, and 1 of 8 paint colors. How many ways can Tommy make a birdhouse? 180 96 56 15

OpenStudy (anonymous):

(I'm correcting my cousins math homework and want to make sure this is right). To solve, it's basically 4+3+8=15 which makes 15 the correct answer, right?

OpenStudy (anonymous):

Or is it\[4*3*8= 96\]?

OpenStudy (anonymous):

Oh wait never mind I remember what I have to do. To everyone who is looking for the answer: It's; \[4*3*8=96\] BECAUSE: The question asks 'AND' which means that he has to choose one of each wood, paint and house styles. IF it was asking 'OR' then it would be:\[4+3+8=15\] BECAUSE: it means that you can only have ONE option/choice.